# LF-Steering: Latent Feature Activation Steering for Enhancing Semantic Consistency in Large Language Models

### v1

@@ 4.1 Datasets and Evaluation Metrics
To ensure a fair comparison, we follow the same evaluation framework used by Yang et al. (2024). This includes using the RobustBOOLQ, RobustSST2, and RobustMRPC datasets for NLU tasks, and the PopQA_Sport and PopQA_capital datasets for NLG tasks. For task performance evaluation, the overall output accuracy across diverse synonymous inputs are employed. To assess the semantic consistency of the LLM, the standard deviation of the overall output accuracy and the mean pairwise cosine similarity are utilized for NLU and NLG datasets respectively.
@@ 4.2 Baseline Methods
In this paper, we compare three baseline methods on the top of our LLM backbone LLama2-7B-Chat. The three model component-level activation steering methods are based on the hidden states of transformer layer and attention heads.
@@
CAA (Rimsky et al., 2024) utilized mean difference layer representations from contrastive samples to steer model behavior [^4] https://github.com/nrimsky/CAA. Following this method, we use the same amount of training data to compute the mass mean steering vector from the top-1 identified transformer layer and the corresponding dataset.
SCS (Yang et al., 2024) introduced an activation steering technique that identifies key components, such as attention heads, that influence semantic consistency. Then, it adjusted the representations of these attention heads to guide the LLM towards greater semantic consistency, thereby improving its overall performance. In this paper, we use the experimental results from their study. Note that, since no specific method name was given to this method in the original work, we refer to it as Semantic Consistency Steering (SCS) for clarity in subsequent experiments.
@@ 4.4 Top-1 Transformer Layer Locating Results
Table 2: Located top k SAE features with manually summarized interpretations and corresponding example activations (highlighted in orange) for the NLU (RobustBOOLQ) and NLG (PopQA_Capital) datasets.
@@
Table 3: Comparison of feature-level and neuron-level LF-Steering methods for enhancing LLM’s semantic consistency.
We utilize the LLama2-7B-Chat model as the target LLM to analyze the most influential layer to the model’s semantic consistency. As shown in Figure 3, we observe a notable trend: the locating accuracy is relatively higher between layers 17 and 32, indicating that the middle to final layers are significantly associated with semantic consistency. Notably, the high locating accuracy in the final layers is evident. It appears these layers may relate to the model’s decision-making process and thus playing a more critical role for steering the final response. For the mid-layers, discrepancies in semantic representations for synonymous inputs may lead to variations in subsequent prediction outcomes. In contrast, the earlier layers primarily capture syntactic and literal information, making it challenging to accurately assess the consistency of their semantic representations; consequently, the locating accuracy is relatively low, hovering around 65%.
4.5 Top-K Features Locating Results
As shown in Table 2, we select several examples to demonstrate the pattern captured by the located SAE features and explain how they may contribute to enhancing the model’s semantic consistency. Specifically, we choose the RobustBOOLQ dataset for NLU and the PopQA_Capital dataset for NLG. We randomly select the top-ranked feature IDs and manually inspect and interpret them using the corresponding datasets. It is found that feature ID 15780 in the RobustBOOLQ dataset is related to “looking for answers”, while feature ID 26247 in the PopQA_Capital dataset is linked to “location-related questions”. Enhancing the activations of these features improves the model’s alignment with varied synonymous expressions, thereby enhancing semantic consistency.
@@ 4.6 Comparing the Effectiveness of Feature-Level and Neuron-Level Steering
To validate the effectiveness of our proposed feature-level steering method compared to neuron-level steering for enhancing LLM’s semantic consistency, we designed a comparative experiment. In specific, LF-Steering (neuron-level) steer the identified top-1 transformer hidden states to enhance LLM’s semantic consistency, whereas LF-Steering (feature-level) steer SAE features proposed in this paper.
@@ 4.7 The Effectiveness of Transformer Layer and SAE Feature Locating
Table 4: Ablation studies on the influence of model layer and SAE features locating strategy. The term “random layer” refers to the strategy of randomly selecting a transformer layer in our steering method, while “random features” refers to randomly selecting SAE features.
@@ 4.8.1 The Impact of Feature Difference Threshold
Figure 4: Performance of our proposed activation steering method across different threshold values on the RobustBOOLQ dataset.
We conduct an analysis of the impact of varying feature difference thresholds in our experimental setup using the RobustBOOLQ dataset. We examine the feature difference threshold $t$ in the set of $\{0.05,0.10,0.15,0.20,0.25,0.30\}$. As shown in Figure 4, the steered model achieves peak accuracy when the $t$ is approximately 0.10. However, we observe a notable decline in predicting accuracy coupled with an increase in standard deviation when the threshold $t$ reaches or exceeds 0.25. This performance decrease can be attributed to an excessively high feature difference threshold, which limits the number of steered features thus results in insufficient steering. These findings emphasize the critical nature of selecting an optimal threshold $t$ for effective steering.
@@ 4.8.2 The Impact of Feature Activation Strength
Figure 5: Performance of our proposed activation steering method across different feature activation values on the RobustBOOLQ dataset.
@@
Figure 5 illustrates that the steered LLM exhibits remarkable robustness across a wide spectrum of $\alpha$ values, ranging from 5.0 to 50.0. Throughout this range, our proposed method consistently enables the backbone LLama2-7B-Chat to improve in both task performance and semantic consistency. However, as anticipated, extreme high value of $\alpha$ (e.g., $\alpha=50.0$) result in performance degradation.
@@ 4.9 Comparison of the Locality of Semantic Consistency Steering Methods
To assess whether our proposed steering method and other steering methods may have negative effects on out-of-domain (OOD) datasets, four datasets were utilized: AG News (Zhang et al., 2015) and IMDB (Maas et al., 2011) for NLU tasks, and CNN/Daily Mail (See et al., 2017) along with XSum (Narayan et al., 2018) for NLG tasks. Accuracy was employed as the evaluation metric for the AG News and IMDB NLU datasets, whereas the ROUGE-L metric (Lin, 2004) was utilized to assess the CNN/Daily Mail and XSum NLG datasets.
As shown in Tables 5 and 6, these steering methods, after enhancing semantic consistency in LLMs, do not negatively impact these OOD datasets, and even lead to improvements on certain datasets. The only exception is ActAdd(Turner et al., 2023), which caused a decrease in classification accuracy for AG News from 70% to 67.20%. This could be due to ActAdd (Turner et al., 2023) using only a single contrastive example for steering, resulting in a larger bias from the single one example. Furthermore, among these steering methods, our proposed LF-Steering method exhibits better locality properties compared to the others. This further highlights the advantage of our proposed feature-level steering approach, which offers more precise steering with minimal interference.
@@ 5 Conclusion
Although our approach achieves significant improvements, we acknowledge that latent representations contributing to semantic inconsistency may span multiple transformer layers in LLMs. Addressing this issue at a single layer may offer only a partial solution, underscoring the importance of a multi-layer activation steering approach. In the furture, we will focus on identifying and steering circuits across multiple transformer layers associated with semantic inconsistency (Elhage et al., 2021; Marks et al., 2024).Moreover, apart from the activation steering approach, it is also possible to analyze the model’s parameters to explicitly identify those parameters that affect semantic consistency and make targeted adjustments. This presents another promising avenue for further investigation.

### v2

@@ 4.1 Datasets and Evaluation Metrics
To ensure a fair comparison, we follow the same evaluation framework used by Yang et al. (2024). This includes using the RobustBOOLQ, RobustSST2, and RobustMRPC datasets for NLU tasks, and the PopQA_Sport and PopQA_capital datasets for NLG tasks. For task performance evaluation, the overall output accuracy across diverse paraphrased inputs are employed. To assess the semantic consistency of the LLM, we use the standard deviation of the overall output accuracy and the mean pairwise cosine similarity for NLU and NLG datasets respectively.
@@ 4.2 Baseline Methods
In this paper, we compare three baseline methods on the top of our LLM backbone LLama2-7B-Chat. The three model component-level activation steering methods are based on the hidden states of transformer layer and attention head outputs.
@@
CAA (Rimsky et al., 2024) utilized mean difference layer representations from contrastive samples to steer model behavior [^4] https://github.com/nrimsky/CAA. Following this method, we use the same amount of training data to compute the mass mean steering vector from the top-1 identified transformer layer and the same corresponding dataset.
SCS (Yang et al., 2024) introduced an activation steering technique that identifies key components, such as attention heads, that influence semantic consistency. Then, it adjusted the representations of these attention heads to guide the LLM towards greater semantic consistency, thereby improving its overall performance. In this paper, we use the experimental results from their study. Note that, since no specific method name was given to this method in the original work, we use Semantic Consistency Steering (SCS) to refer to this method for clarity in subsequent experiments.
@@ 4.4 Top-1 Transformer Layer Locating Results
Table 2: Located key SAE features with manually summarized interpretations and corresponding example activations (highlighted in orange) for the NLU (RobustBOOLQ) and NLG (PopQA_Capital) datasets.
@@
Table 3: Comparison of feature-level and neuron-level LF-Steering methods for enhancing LLM’s semantic consistency and task performance.
We use the LLama2-7B-Chat model and the top-1 transformer layer locating method mentioned in Section 3.1 to analyze the contributions of transformer layers to the model’s semantic consistency. As shown in Figure 3, we observe a notable trend: the locating accuracy is relatively higher between layers 17 and 32, indicating that the middle to final layers are significantly associated with semantic consistency. Notably, the high locating accuracy in the final layers is evident. It appears these layers may relate to the model’s decision-making process and thus playing a more critical role for steering the final response. For the mid-layers, discrepancies in semantic representations for paraphrased inputs may lead to variations in subsequent prediction outcomes. In contrast, the earlier layers primarily capture syntactic and literal information (DBLP:conf/acl/LiuKL024), making it challenging to accurately assess the consistency of their semantic representations; consequently, the locating accuracy is relatively low, hovering around 65%.
4.5 Interpretation of Feature Representations
As shown in Table 2, we select several examples to demonstrate the pattern captured by the located key SAE features and explain how they may contribute to enhancing the model’s semantic consistency. Specifically, we choose the RobustBOOLQ dataset for NLU and the PopQA_Capital dataset for NLG. We randomly select the top-ranked feature IDs and manually inspect and interpret them using the corresponding datasets. It is found that feature ID 15780 in the RobustBOOLQ dataset is related to “looking for answers”, while feature ID 26247 in the PopQA_Capital dataset is linked to “location-related questions”. Enhancing the activations of these features improves the model’s alignment with various semantically equivalent expressions, while exerting lower interference on other features, thereby resulting in improved performance in enhancing semantic consistency.
@@ 4.6 Comparing the Effectiveness of Feature-Level and Neuron-Level Steering
To validate the effectiveness of our proposed feature-level steering method compared to neuron-level steering for enhancing LLM’s semantic consistency, we designed a comparative experiment. In specific, LF-Steering (neuron-level) steer the identified top-1 transformer hidden states to enhance LLM’s semantic consistency, whereas LF-Steering (feature-level) steer the SAE features.
@@ 4.7 The Effectiveness of Transformer Layer and SAE Feature Locating
Table 4: Ablation studies on the influence of model layer and SAE features locating strategy. The term “random layer” refers to randomly selecting a transformer layer, while “random features” refers to randomly selecting SAE features.
@@ 4.8.1 The Impact of Feature Difference Threshold
Figure 4: Performance of our proposed activation steering method across different threshold values.
We conduct an analysis of the impact of varying feature difference thresholds in our experimental setup using the RobustBOOLQ dataset. We examine the feature difference threshold $t$ in the set of $\{0.05,0.10,0.15,0.20,0.25,0.30\}$. As shown in Figure 4, the steered model achieves peak accuracy when the $t$ is approximately 0.10. However, we observe a notable decline in predicting accuracy coupled with an increase in standard deviation when the threshold $t$ reaches or exceeds 0.25. This performance decrease can be attributed to an excessively high feature difference threshold, which limits the number of steered features thus results in insufficient steering. These findings emphasize the critical nature of selecting an optimal threshold $t$ for effective steering. In this paper, we set $t$ to a value of $0.1$.
@@ 4.8.2 The Impact of Feature Activation Strength
Figure 5: Performance of our proposed activation steering method across different feature activation values.
@@
Figure 5 illustrates that the steered LLM exhibits remarkable robustness across a wide spectrum of $\alpha$ values, ranging from 5.0 to 50.0. Throughout this range, our proposed method consistently enables the backbone LLama2-7B-Chat to improve in both task performance and semantic consistency. However, as anticipated, extreme high value of $\alpha$ (e.g., $\alpha=50.0$) result in performance degradation. In this paper, we set $\alpha$ to a value of $20$.
@@ 4.9 Comparison of the Locality of Semantic Consistency Steering Methods
To assess the locality (gu2024model) of our proposed steering method and other steering methods, we tested whether these methods might have negative effects on out-of-domain (OOD) datasets. In specific, four OOD datasets were utilized: AG News (Zhang et al., 2015) and IMDB (Maas et al., 2011) for NLU tasks, and CNN/Daily Mail (See et al., 2017) along with XSum (Narayan et al., 2018) for NLG tasks. Accuracy was employed as the evaluation metric for the AG News and IMDB NLU datasets, whereas the ROUGE-L metric (Lin, 2004) was utilized to assess the CNN/Daily Mail and XSum NLG datasets.
As shown in Tables 5 and 6, these steering methods, after enhancing semantic consistency in LLMs, do not negatively impact these OOD datasets, and even lead to improvements on certain datasets. The only exception is ActAdd (Turner et al., 2023), which caused a decrease in classification accuracy for AG News from 70% to 67.20%. This could be due to ActAdd (Turner et al., 2023) using only a single contrastive example for steering, resulting in a larger bias from the single one example. Furthermore, among these steering methods, our proposed LF-Steering method exhibits better locality properties compared to the others. This further highlights the advantage of our proposed feature-level steering approach, which offers more precise steering with lower interference.
@@ 5 Conclusion
Although our approach achieves significant improvements, we acknowledge that latent representations contributing to semantic inconsistency may span multiple transformer layers in LLMs. Addressing this issue at a single layer may offer only a partial solution, underscoring the importance of a multi-layer activation steering approach. In the furture, we will focus on identifying and steering circuits across multiple transformer layers to enhance the LLM’s semantic consistency (Elhage et al., 2021; Marks et al., 2024).